# Stability of Ecological SystemsCan Chen, Xu-Wen Wang, and Yang-Yu Liu

### 12 Dec 2023 v1

@@ 5.1 Lyapunov stability of the GLV model
Before talking about the stability of the GLV model, we first introduce the notion of Persidskii-type systems. Persidskii-type systems are typical examples that admit diagonal-type Lyapunov functions [85, 86, 194, 201]. We need to introduce a few concepts to define Persidskii-type systems. A function $\textbf{f}:\mathbb{R}^{S}\to\mathbb{R}^{S}:\textbf{x}\mapsto\textbf{f}(\textbf{x})$ is called diagonal if, the $i$th component of f, i.e., $f_{i}$, is a function of $x_{i}$ alone. A function $f:\mathbb{R}\to\mathbb{R}$ is said to be in sector $[l,u]$ if $\forall x\in\mathbb{R}$, $f(x)$ lies between $lx$ and $ux$, i.e.,
@@
Using the above diagonal-type Lyapunov function, Goh first proved that for the GLV model (10), if the interaction matrix A is diagonally stable, then the non-trivial equilibrium point $\textbf{x}^{*}$ in the positive orthant is globally asymptotically stable (in the sense of Lyapunov) [107]. Therefore, diagonal stability allows the GLV model to lie in the unique fixed-point phase, as depicted by Bunin [57]. Here, we provide an example of a 3-by-3 diagonally stable matrix and the corresponding vector field plot of the GLV model, see Fig. 8. Clearly, the non-trivial equilibrium point in the positive orthant is globally asymptotically stable.
@@
The diagonal stability analysis has been applied to analyze the stability of various GLV-derived models. Following Goh’s findings, Wörz-Busekros offered a sufficient condition for the global stability of the GLV model with continuous-time delay via diagonal stability analysis [266]. Later on, Beretta and Takeuchi considered the global asymptotic stability of diffusion models with multiple species heterogeneous patches, in which each patch is governed by the GLV model with continuous-time delay [43]. Concurrently, Beretta and Solimano generalized the Wörz-Busekros’s outcome by considering a non-negative linear vector function of the species [42]. In addition, Kon exploited the diagonal stability theory to determine the stability of the GLV model with an age structure [142]. The notion of diagonal stability was further applied to more general ecological models such as Kolmogorov systems [124], where the GLV model is encompassed as a specific instance. Significantly, the stability properties of many quasi-polynomial dynamical systems, often used to represent many biochemical processes, can be studied through an equivalent GLV model that has a much simpler form [120, 171]. This is based on the fact that the former can be transformed into the latter with some appropriate changes of variables.
@@ 5.2 Characterizations of diagonally stable matrices
Grilli et al. imposed a condition of negative definiteness on A (i.e., all the eigenvalues of $\textbf{A}+\textbf{A}^{\top}$ are negative) to ensure the diagonal stability of A for studying the feasibility and coexistence of large random ecological systems [110]. It is known that a negative definite matrix is also diagonally stable, and the condition is much easier to verify and characterized for random matrices. Later on, Gibbs et al. [100] estimated the “rightmost” eigenvalue of $(\textbf{A}+\textbf{A}^{\top})/2$ with random interactions and proved that if
@@
Recently, necessary and sufficient diagonal stability conditions for matrices associated with special interconnection or network structures were studied [19, 20, 21, 22, 233, 262]. If an ecological system described by the GLV model exhibits these network structures, it will be effective and efficient to determine its Lyapunov stability through the diagonal stability of the interaction matrix A. We review these network structures as follows.
@@ 5.2.1 Negative feedback cyclic structure
In the negative feedback cyclic structure, the intermediate species play a facilitating role for the subsequent species, while the final species exerts inhibitory effects on the initial species, see Fig. 9. It has been shown that the interaction matrix A is Hurwitz, i.e., $\text{Re}[\lambda(\textbf{A})]<0$, if it satisfies the so-called secant criterion [251, 255, 265], i.e.,
@@
When $\alpha_{i}$ are equal, the secant criterion (22) is also necessary for A to be Hurwitz. Surprisingly, this secant criterion derived for linear stability is also a necessary and sufficient condition for diagonal stability of the corresponding class of matrices [21].
Note that a simple necessary condition for A to be diagonal stable is that all the diagonal elements $\textbf{A}_{ii}$ be negative. Since scaling the rows of A by positive constants does not change its diagonal stability, one can safely assume $\textbf{A}_{ii}=-1$. Moreover, a reducible matrix A can always be transformed into an upper-triangle form with a suitable permutation, and permutation does not change diagonal stability. Considering these two points, the secant criterion can be further generalized as follows [19]. Any matrix A that can be transformed via a suitable permutation P to the form of
@@ 5.2.2 Cactus structure
Arcak further generalized the secant criterion to multiple cycles when the weighted digraph of the interaction matrix A, denoted as $G(\textbf{A})$, possesses a “cactus” structure, i.e., any pair of distinct simple cycles have at most one common node [19]. Apparently, the negative feedback cyclic structure corresponds to a single cycle and is just a special case of the cactus structure. The digraph $G(\textbf{A})$ is defined to represent the off-diagonal entries of A that has $S$ nodes and there is a directed edge $(j\to i)$ with weight $\textbf{A}_{ij}$ if and only if $\textbf{A}_{ij}\neq 0$. Self-loops corresponding to the diagonal entries $\textbf{A}_{ii}$ are excluded from $G(\textbf{A})$.
Arcak made two assumptions about A without loss of generality: (i) $\textbf{A}_{ii}=-1$ for $i=1,\dots,S$; (ii) $G(\textbf{A})$ is strongly connected, or equivalently, A is irreducible. Arcak then defined an undirected graph for describing which cycles of the digraph $G(\textbf{A})$ intersect. Subsequently, he constructed a spanning tree in the undirected graph and simultaneously assigned directions to its edges to create an arborescence, i.e., a digraph in which every node can be reached from the root by one and only one path, see Fig. 10a, b, c. Using the hierarchy of the cycles of $G(\textbf{A})$ established by this arborescence presentation, he sequentially generated dinequalities for the gains of each simple cycle.
Figure 10: Cactus and connected circle structures. (a) Digraph $G(\textbf{A})$ corresponding to the matrix A with simple cycles $\{1,4,1\},\{1,2,3,1\},\{1,7,8,1\}$, and $\{4,5,6,4\}$ labeled as 1, 2, 3 and 4, respectively. Note that any pair of the cycles have at most one common node. (b) Undirected graph that describes which cycles in $G(\textbf{A})$ intersect. (c) Arborescence constructed on the undirected graph in (b) according to the broadcasting algorithm with cycle 1 selected as the root (shown in cyan). (d) Digraph with the connected circle structure where any pair of the cycles have at most one common edge or one common node. Panels a, b, c were redrawn from [19].
@@
Arcak then outlined a systematic procedure for constructing Lyapunov functions based on the above stability condition. Notably, he illustrated the procedure with the GLV model for ecological systems [19].
Later on, Wang and Nešić extended the small gain condition to a more general circle structure called connected circles, where each pair of distinct simple cycles have at most one common edge or a common node [262], see Fig. 10d. In addition to the two assumptions (i) and (ii) made by Arcak, the authors further assumed that (iii) A has non-negative off-diagonal elements. Define $\mathcal{L}_{j}=\{e_{j}^{(1)},e_{j}^{(2)},\dots,e_{j}^{(n_{j}-1)}\}$ as the set of edges traversed by cycle $j$. Then the stability condition can be modified as follows: the interaction matrix A satisfying the above three assumptions (i-iii) is diagonally stable if and only if there exist constants $\theta_{j}^{(i)}>0$ such that
@@ 5.2.3 Rank-one structure
Recently, Simpson-Porco and Monshizadeh explored necessary and sufficient conditions for the diagonal stability of A with a rank-one network structure [233] defined as
@@ 6 D-stability analysis
The notion of D-stability was originally introduced by Arrow and McManus [24] and Enthoven and Arrow [90] in the late 1950s. D-stability can be considered as a weaker version of diagonal stability, see Fig. 11. A matrix A is called D-stable if for any positive diagonal matrix X, the matrix XA is stable (in terms of $\text{Re}[\lambda(\textbf{X}\textbf{A})]\leq 0$). Clearly, the definition of D-stability is particularly relevant to the GLV model, where its community matrix is represented as $\textbf{M}=\textbf{X}\textbf{A}$. In this section, we first investigate the characterizations of D-stable matrices and then discuss existing results regarding D-stability in the context of the GLV model.
@@ 6.1 Characterizations of D-stable matrices
Explicitly characterizing D-stable matrices is only known for dimension less than or equal to four [159]. For example, a matrix $-\textbf{A}\in\mathbb{R}^{4\times 4}$ is D-stable if and only if all the principal minors of A are positive and $f(x,y,z)>0$ for all positive $x$, $y$, and $z$, where
@@
and $\textbf{D}=\text{diag}(1,x,y,z)$. Here, $E_{i}(\cdot)$ denotes the sums of the order $i$ principal minors of a given matrix. However, the problem becomes highly challenging as the dimension grows larger [50, 100, 159, 210]. Nonetheless, numerous sufficient conditions have been proposed [132, 151, 189]. Some of the better-known ones are:
  • If A is diagonally stable, then it is D-stable [24]. In fact, it is essentially the condition that Arrow and McManus offered.
@@
  • If A is triangular with $\textbf{A}_{ii}<0$ for $i=1,2,\dots,n$, then it is D-stable [132]. This is the most straightforward condition for D-stability.
  • If A is sign stable, then it is D-stable [132].
  • The element-wise product of P and A is stable for each positive definite symmetric matrix P [208].
Regrettably, none of these conditions is necessary for D-stability. More sufficient conditions can be found in [132].
Figure 11: Schematic containment plot of matrix stability notions. Detailed proofs of the relations among these stability notions can be found in [159]. This figure was redrawn from [222].
@@ 6.2 Results with the GLV model
Here, we present a series of findings and speculation regarding D-stability in the context of analyzing the GLV model. Chu offered a solvable Lie algebraic condition for the equivalence of four stability notions – linear stability, D-stability, total stability (defined in Section 7), and diagonal stability for the GLV model [72]. Given a Lie algebra $\mathbb{L}$, let’s define the inductive sequence as
@@
The Lie algebra $\mathbb{L}$ is referred to as solvable if there exists a positive integer $k>0$ such that $\mathbb{L}^{(k)}=\{\textbf{0}\}$. The author proved that given the GLV model (10), if the two matrices, defined as $(\textbf{A}+\textbf{A}^{\top})/2$ and $(\textbf{A}-\textbf{A}^{\top})/2$, generate a solvable Lie algebra, then the linear stability, D-stability, total stability, and diagonal stability of the interaction matrix A are equivalent. Moreover, although diagonal stability is not equivalent to Lyapunov stability in general, based on intensive numerical simulations, Rohr et al. conjectured that for mutualistic ecological systems captured by the GLV model, if the interaction matrix A is stable (in the sense of Lyapunov), then A is D-stable [213]. This conjecture has an important consequence for modeling mutualistic ecological systems with the GLV model, implying that if A is stable (in the sense of Lyapunov), any feasible equilibrium point is globally stable, corresponding to the unique fixed-point phase, as illustrated by Bunin [57]. Similarly, Gibbs et al. observed that large random matrices are D-stable almost surely (i.e., the set of positive diagonal matrices leading to instability has measure zero), so feasible unstable equilibrium points are very unlikely for the GLV model with random interactions [100].
@@ 7 Total Stability
The notion of D-stability is also closely linked to another stability notion termed total stability. A matrix A is called totally stable if any principal sub-matrix of A is D-stable [121, 206]. Any principal sub-matrix of a totally stable matrix is also totally stable. In addition, total stability is a stronger notion compared to D-stability, but is weaker than diagonal stability, as illustrated in Fig. 11. Consequently, any totally stable matrix is inherently D-stable, and the class of totally stable matrices is closed under transposition and multiplication by a positive diagonal matrix. This implies that total stability holds simultaneously for both the interaction matrix A and the community matrix M of the GLV model [157]. A direct way to characterize a totally stable matrix is to check the D-stability of all its principal sub-matrices, which is computationally expensive. However, explicit characterizations are only known for $S\leq 3$ [156, 157], similar to diagonal stability and D-stability. A necessary condition of total stability is that all the principal minors of A of odd orders are negative, while those of even orders are positive [157]. Here, we provide an example of a 3-by-3 totally stable matrix that is not diagonally stable as follows (borrowed from [157]):
@@
Ecologically, the principal sub-matrix of the interaction matrix refers to the interactions between the subset of species that remains after removal or extinction of certain species. Therefore, total stability is related to the so-called “species-deletion stability” concept introduced in [200], implying that the property is preserved after the removal or extinction of any group of species from the initial composition [159]. In the context of the GLV model, the total stability of the interaction matrix suggests that the remaining species will reach a new locally stable equilibrium after the removal or extinction of certain species.

### 07 Feb 2024 v2

@@ 5.1 Lyapunov stability of the GLV model
Before talking about the stability of the GLV model, we first introduce the notion of Persidskii-type systems. Persidskii-type systems are typical examples that admit diagonal-type Lyapunov functions efimov2021analysis ; efimov2019robust ; Persidskii-ARC-69 ; platonov2023asymptotic . We need to introduce a few concepts to define Persidskii-type systems. A function $\textbf{f}:\mathbb{R}^{S}\to\mathbb{R}^{S}:\textbf{x}\mapsto\textbf{f}(\textbf{x})$ is called diagonal if, the $i$th component of f, i.e., $f_{i}$, is a function of $x_{i}$ alone. A function $f:\mathbb{R}\to\mathbb{R}$ is said to be in sector $[l,u]$ if $\forall x\in\mathbb{R}$, $f(x)$ lies between $lx$ and $ux$, i.e.,
@@
Using the above diagonal-type Lyapunov function, Goh first proved that for the GLV model (10), if the interaction matrix A is diagonally stable, then the non-trivial equilibrium point $\textbf{x}^{*}$ in the positive orthant is globally asymptotically stable (in the sense of Lyapunov) goh1977global . Therefore, diagonal stability allows the GLV model to lie in the unique fixed-point phase, as depicted by Bunin bunin2017ecological . Here, we provide an example of a 3-by-3 diagonally stable matrix and the corresponding vector field plot of the GLV model, see Figure 8. Clearly, the non-trivial equilibrium point in the positive orthant is globally asymptotically stable.
@@
The diagonal stability analysis has been applied to analyze the stability of various GLV-derived models. Following Goh’s findings, Wörz-Busekros offered a sufficient condition for the global stability of the GLV model with continuous-time delay via diagonal stability analysis worz1978global . Later on, Beretta and Takeuchi considered the global asymptotic stability of diffusion models with multiple species heterogeneous patches, in which each patch is governed by the GLV model with continuous-time delay beretta1988global . Concurrently, Beretta and Solimano generalized the Wörz-Busekros’s outcome by considering a non-negative linear vector function of the species beretta1988generalization . In addition, Kon exploited the diagonal stability theory to determine the stability of the GLV model with an age structure kon2011age . The notion of diagonal stability was further applied to more general ecological models such as Kolmogorov systems hou2015global , where the GLV model is encompassed as a specific instance. Significantly, the stability properties of many quasi-polynomial dynamical systems, often used to represent many biochemical processes, can be studied through an equivalent GLV model that has a much simpler form hernandez1997lotka ; motee2012stability . This is based on the fact that the former can be transformed into the latter with some appropriate changes of variables.
@@ 5.2 Characterizations of diagonally stable matrices
Grilli et al. imposed a condition of negative definiteness on A (i.e., all the eigenvalues of $\textbf{A}+\textbf{A}^{\top}$ are negative) to ensure the diagonal stability of A for studying the feasibility and coexistence of large random ecological systems grilli2017feasibility . It is known that a negative definite matrix is also diagonally stable, and the condition is much easier to verify and characterized for random matrices. Later on, Gibbs et al. gibbs2018effect estimated the “rightmost” eigenvalue of $(\textbf{A}+\textbf{A}^{\top})/2$ with random interactions and proved that if
@@
Recently, necessary and sufficient diagonal stability conditions for matrices associated with special interconnection or network structures were studied arcak2011diagonal ; arcak2008passivity ; arcak2006diagonal ; 4434115 ; simpson2021diagonal ; wang2012diagonal . If an ecological system described by the GLV model exhibits these network structures, it will be effective and efficient to determine its Lyapunov stability through the diagonal stability of the interaction matrix A. We review these network structures as follows.
@@ 5.2.1 Negative feedback cyclic structure
In the negative feedback cyclic structure, the intermediate species play a facilitating role for the subsequent species, while the final species exerts inhibitory effects on the initial species, see Figure 9. It has been shown that the interaction matrix A is Hurwitz, i.e., $\text{Re}[\lambda(\textbf{A})]<0$, if it satisfies the so-called secant criterion Thron-BMB-91 ; Tyson-Book-78 ; wimmer2009diagonal , i.e.,
@@
When $\alpha_{i}$ are equal, the secant criterion (22) is also necessary for A to be Hurwitz. Surprisingly, this secant criterion derived for linear stability is also a necessary and sufficient condition for diagonal stability of the corresponding class of matrices arcak2006diagonal .
Note that a simple necessary condition for A to be diagonal stable is that all the diagonal elements $\textbf{A}_{ii}$ be negative. Since scaling the rows of A by positive constants does not change its diagonal stability, one can safely assume $\textbf{A}_{ii}=-1$. Moreover, a reducible matrix A can always be transformed into an upper-triangle form with a suitable permutation, and permutation does not change diagonal stability. Considering these two points, the secant criterion can be further generalized as follows arcak2011diagonal . Any matrix A that can be transformed via a suitable permutation P to the form of
@@ 5.2.2 Cactus structure
Arcak further generalized the secant criterion to multiple cycles when the weighted digraph of the interaction matrix A, denoted as $G(\textbf{A})$, possesses a “cactus” structure, i.e., any pair of distinct simple cycles have at most one common node arcak2011diagonal . Apparently, the negative feedback cyclic structure corresponds to a single cycle and is just a special case of the cactus structure. The digraph $G(\textbf{A})$ is defined to represent the off-diagonal entries of A that has $S$ nodes and there is a directed edge $(j\to i)$ with weight $\textbf{A}_{ij}$ if and only if $\textbf{A}_{ij}\neq 0$. Self-loops corresponding to the diagonal entries $\textbf{A}_{ii}$ are excluded from $G(\textbf{A})$.
Arcak made two assumptions about A without loss of generality: (i) $\textbf{A}_{ii}=-1$ for $i=1,\dots,S$; (ii) $G(\textbf{A})$ is strongly connected, or equivalently, A is irreducible. Arcak then defined an undirected graph for describing which cycles of the digraph $G(\textbf{A})$ intersect. Subsequently, he constructed a spanning tree in the undirected graph and simultaneously assigned directions to its edges to create an arborescence, i.e., a digraph in which every node can be reached from the root by one and only one path, see Figure 10a, b, c. Using the hierarchy of the cycles of $G(\textbf{A})$ established by this arborescence presentation, he sequentially generated dinequalities for the gains of each simple cycle.
Figure 10: Cactus and connected circle structures. (a) Digraph $G(\textbf{A})$ corresponding to the matrix A with simple cycles $\{1,4,1\},\{1,2,3,1\},\{1,7,8,1\}$, and $\{4,5,6,4\}$ labeled as 1, 2, 3 and 4, respectively. Note that any pair of the cycles have at most one common node. (b) Undirected graph that describes which cycles in $G(\textbf{A})$ intersect. (c) Arborescence constructed on the undirected graph in (b) according to the broadcasting algorithm with cycle 1 selected as the root (shown in cyan). (d) Digraph with the connected circle structure where any pair of the cycles have at most one common edge or one common node. Panels a, b, c were redrawn from arcak2011diagonal .
@@
Arcak then outlined a systematic procedure for constructing Lyapunov functions based on the above stability condition. Notably, he illustrated the procedure with the GLV model for ecological systems arcak2011diagonal .
Later on, Wang and Nešić extended the small gain condition to a more general circle structure called connected circles, where each pair of distinct simple cycles have at most one common edge or a common node wang2012diagonal , see Figure 10d. In addition to the two assumptions (i) and (ii) made by Arcak, the authors further assumed that (iii) A has non-negative off-diagonal elements. Define $\mathcal{L}_{j}=\{e_{j}^{(1)},e_{j}^{(2)},\dots,e_{j}^{(n_{j}-1)}\}$ as the set of edges traversed by cycle $j$. Then the stability condition can be modified as follows: the interaction matrix A satisfying the above three assumptions (i-iii) is diagonally stable if and only if there exist constants $\theta_{j}^{(i)}>0$ such that
@@ 5.2.3 Rank-one structure
Recently, Simpson-Porco and Monshizadeh explored necessary and sufficient conditions for the diagonal stability of A with a rank-one network structure simpson2021diagonal defined as
@@ 6 D-stability analysis
The notion of D-stability was originally introduced by Arrow and McManus arrow1958note and Enthoven and Arrow enthoven1956theorem in the late 1950s. D-stability can be considered as a weaker version of diagonal stability, see Figure 11. A matrix A is called D-stable if for any positive diagonal matrix X, the matrix XA is stable (in terms of $\text{Re}[\lambda(\textbf{X}\textbf{A})]\leq 0$). Clearly, the definition of D-stability is particularly relevant to the GLV model, where its community matrix is represented as $\textbf{M}=\textbf{X}\textbf{A}$. In this section, we first investigate the characterizations of D-stable matrices and then discuss existing results regarding D-stability in the context of the GLV model.
@@ 6.1 Characterizations of D-stable matrices
Explicitly characterizing D-stable matrices is only known for dimension less than or equal to four logofet2005stronger . For example, a matrix $-\textbf{A}\in\mathbb{R}^{4\times 4}$ is D-stable if and only if all the principal minors of A are positive and $f(x,y,z)>0$ for all positive $x$, $y$, and $z$, where
@@
and $\textbf{D}=\text{diag}(1,x,y,z)$. Here, $E_{i}(\cdot)$ denotes the sums of the order $i$ principal minors of a given matrix. However, the problem becomes highly challenging as the dimension grows larger bhaya1993discrete ; gibbs2018effect ; logofet2005stronger ; Redheffer-SIAM-85 . Nonetheless, numerous sufficient conditions have been proposed johnson1974sufficient ; leite2003improved ; peaucelle2000new . Some of the better-known ones are:
  • 1.
    If A is diagonally stable, then it is D-stable arrow1958note . In fact, it is essentially the condition that Arrow and McManus offered.
  • 2.
@@
  • 4.
    If A is triangular with $\textbf{A}_{ii}<0$ for $i=1,2,\dots,n$, then it is D-stable johnson1974sufficient . This is the most straightforward condition for D-stability.
  • 5.
    If A is sign stable, then it is D-stable johnson1974sufficient .
  • 6.
    The element-wise product of P and A is stable for each positive definite symmetric matrix P r1974hadamard .
Regrettably, none of these conditions is necessary for D-stability. More sufficient conditions can be found in johnson1974sufficient .
Figure 11: Schematic containment plot of matrix stability notions. Detailed proofs of the relations among these stability notions can be found in logofet2005stronger . This figure was redrawn from saeedian2022effect .
@@ 6.2 Results with the GLV model
Here, we present a series of findings and speculation regarding D-stability in the context of analyzing the GLV model. Chu offered a solvable Lie algebraic condition for the equivalence of four stability notions – linear stability, D-stability, total stability (defined in Section 7), and diagonal stability for the GLV model chu2007equivalent . Given a Lie algebra $\mathbb{L}$, let’s define the inductive sequence as
@@
The Lie algebra $\mathbb{L}$ is referred to as solvable if there exists a positive integer $k>0$ such that $\mathbb{L}^{(k)}=\{\textbf{0}\}$. The author proved that given the GLV model (10), if the two matrices, defined as $(\textbf{A}+\textbf{A}^{\top})/2$ and $(\textbf{A}-\textbf{A}^{\top})/2$, generate a solvable Lie algebra, then the linear stability, D-stability, total stability, and diagonal stability of the interaction matrix A are equivalent. Moreover, although diagonal stability is not equivalent to Lyapunov stability in general, based on intensive numerical simulations, Rohr et al. conjectured that for mutualistic ecological systems captured by the GLV model, if the interaction matrix A is stable (in the sense of Lyapunov), then A is D-stable rohr2014structural . This conjecture has an important consequence for modeling mutualistic ecological systems with the GLV model, implying that if A is stable (in the sense of Lyapunov), any feasible equilibrium point is globally stable, corresponding to the unique fixed-point phase, as illustrated by Bunin bunin2017ecological . Similarly, Gibbs et al. observed that large random matrices are D-stable almost surely (i.e., the set of positive diagonal matrices leading to instability has measure zero), so feasible unstable equilibrium points are very unlikely for the GLV model with random interactions gibbs2018effect .
@@ 7 Total Stability
The notion of D-stability is also closely linked to another stability notion termed total stability. A matrix A is called totally stable if any principal sub-matrix of A is D-stable hofbauer1998evolutionary ; quirk1965qualitative . Any principal sub-matrix of a totally stable matrix is also totally stable. In addition, total stability is a stronger notion compared to D-stability, but is weaker than diagonal stability, as illustrated in Figure 11. Consequently, any totally stable matrix is inherently D-stable, and the class of totally stable matrices is closed under transposition and multiplication by a positive diagonal matrix. This implies that total stability holds simultaneously for both the interaction matrix A and the community matrix M of the GLV model logofet2018matrices . A direct way to characterize a totally stable matrix is to check the D-stability of all its principal sub-matrices, which is computationally expensive. However, explicit characterizations are only known for $S\leq 3$ logofet2018matrices ; logofet1987on , similar to diagonal stability and D-stability. A necessary condition of total stability is that all the principal minors of A of odd orders are negative, while those of even orders are positive logofet2018matrices . Here, we provide an example of a 3-by-3 totally stable matrix that is not diagonally stable as follows (borrowed from logofet2018matrices ):
@@
Ecologically, the principal sub-matrix of the interaction matrix refers to the interactions between the subset of species that remains after removal or extinction of certain species. Therefore, total stability is related to the so-called “species-deletion stability” concept introduced in pimm1982food , implying that the property is preserved after the removal or extinction of any group of species from the initial composition logofet2005stronger . In the context of the GLV model, the total stability of the interaction matrix suggests that the remaining species will reach a new locally stable equilibrium after the removal or extinction of certain species.
8 Sector stability analysis
The notion of sector stability, proposed by Goh, focuses on the stability properties of semi-feasible equilibrium points (i.e., $x_{i}^{*}=0$ for some $i$) for studying interesting processes in ecology, e.g., succession and extinction case1979global ; goh1978sector . A semi-feasible equilibrium point is called (locally) sector stable if every trajectory of the system that starts within a non-negative neighborhood remains in the same or an even larger non-negative neighborhood and eventually converges to that equilibrium point. The definition is analogous to that of (local) asymptotic stability. However, sector stability restricts the trajectories within a non-negative part of an open neighborhood of the equilibrium point.
Let’s consider a generic population dynamics model defined as
$\dot{x}_{i}=x_{i}f_{i}(x_{1},x_{2},\dots,x_{S})$ (26)
for $i=1,2,\dots,S$, where $f_{i}$ have continuous partial derivatives at every finite point in the state space. Let $I_{0}=\{i\text{ }|\text{ }x_{i}^{*}=0\}$ and $I_{1}=\{i\text{ }|\text{ }x_{i}^{*}>0\}$. Goh proved the semi-feasible equilibrium point $\textbf{x}^{*}$ of the generic population model (26) is locally sector stable if all the eigenvalues of the community matrix defined as $\textbf{M}_{ij}=x_{i}^{*}\frac{\partial f_{i}}{\partial x_{j}}|_{\textbf{x}=\textbf{x}^{*}}$ have negative real parts and $f_{i}(\textbf{x}^{*})<0$ for all $i\in I_{0}$ goh1978sector . Global sector stability results were also established by Goh through Lyapunov theory. Define $U=\{\textbf{x}\text{ }|\textbf{ }x_{i}>0\text{ for }i\in I_{1}\text{ and }x_{i}\geq 0\text{ for }i\in I_{0}\}$. A semi-feasible equilibrium point is called globally sector stable if it is sector stable relative to the set $U$. The semi-feasible equilibrium point $\textbf{x}^{*}$ of the generic population model (26) is globally sector stable if there exists positive constants $d_{1},d_{2},\dots,d_{S}$ such that at every point in $U$, the function
$V(\textbf{x})=\sum_{i=1}^{S}d_{i}(x_{i}-x_{i}^{*})f_{i}(\textbf{x})\leq 0,$
and it does not vanish identically along a solution of (26) except for $\textbf{x}=\textbf{x}^{*}$ goh1978sector . The result establishes valuable conditions for global sector stability in the GLV model (10): (i) there exists a positive diagonal matrix $\textbf{D}=\text{diag}(d_{1},d_{2},\dots,d_{S})$ such that $\textbf{A}+\textbf{A}^{\top}\textbf{D}$ D is negative semi-definite (i.e., A is diagonally semi-stable); (ii) the expressions $r_{i}+\sum_{j=1}^{S}\textbf{A}_{ij}x_{j}^{*}\leq 0$ for all $i\in I_{0}$; (iii) the function $\frac{1}{2}(\textbf{x}-\textbf{x}^{*})^{\top}(\textbf{D}\textbf{A}+\textbf{A}^{\top}\textbf{D})(\textbf{x}-\textbf{x}^{*})+\sum_{i\in I_{0}}d_{i}x_{i}(r_{i}+\sum_{j=1}^{S}\textbf{A}_{ij}x_{j}^{*})$ does not vanish identically along any solution of (10) except for $\textbf{x}=\textbf{x}^{*}$.
However, the above global sector stability condition is difficult to verify in practice (especially when $S>2$). Thus, Goh came up with a conservative but simpler result. Suppose that there exists a constant matrix G such that
$\begin{split}&\frac{\partial f_{i}(\textbf{x})}{\partial x_{i}}|_{\textbf{x}=\textbf{x}^{*}}\leq\textbf{G}_{ii}<0\quad\text{for }i=1,2,\dots,S,\\ &\Big{|}\frac{\partial f_{i}(\textbf{x})}{\partial x_{j}}|_{\textbf{x}=\textbf{x}^{*}}\Big{|}\leq\textbf{G}_{ij}\quad\text{for }i\neq j\end{split}$ (27)
[…]
$\begin{cases}\dot{x}_{1}&=x_{1}(11.7-4x_{1}-0.2x_{2}-0.1x_{3})\\ \dot{x}_{2}&=x_{2}(1.2-0.8x_{1}-x_{2}-0.2x_{3})\\ \dot{x}_{3}&=x_{3}(3-2x_{1}-x_{2}-2x_{3})\end{cases}.$
[…]
Figure 12: Trajectories of the GLV model with three initial conditions closely positioned to the other three semi-feasible equilibrium points. The globally sector stable semi-feasible equilibrium point is marked by the red star.
The feasibility requirement for equilibrium points in complex ecological models such as the GLV model drastically restricts their parameter space, especially when assuming random interactions. Consequently, most equilibrium points are semi-feasible, which highlights the importance and necessity of sector stability. Furthermore, Goh’s results indicate that a semi-feasible equilibrium point will likely be globally sector stable if the strength of self-regulating interactions surpasses that of interspecific interactions goh1978sector .